# Learning in Structured Stackelberg Games

### v2

@@ 4.1 Learning without Contexts
Next, we present the definition of pseudo-dimension [34], which is complexity measure that characterizes a hypothesis class and in our setting helps us reason about how difficult the learning task of the leader is. Informally speaking, the more complex the hypothesis class, $\pazocal{P}$, the harder it is for the leader to distinguish between strategies, $\mathbf{x}$, which parameterize the class, and therefore, more samples are needed to learn.
Definition 4.4 (Pseudo-dimension).
Using Lemma˜4.3 and results from the data-driven algorithm design literature [6, 12, 14], we immediately get the following bound on the pseudo-dimension.
Theorem 4.5 (Adaptation of Lemma 3.10 in Balcan et al. [9]).
It is known that the finiteness of the pseudo-dimension of a function class is a sufficient condition to show a generalization guarantee for the class [34, 24]. In our case, the above bound implies a generalization guarantee with function
Discussion.
We note that the above generalization result is particularly strong. Informally, it asserts that for any distribution over the follower types, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far. This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [29].
4.2 Learning with Contexts: Problem Setting and Definitions
In this section, we introduce the study of the statistical complexity of distributional learning in structured Stackelberg games. The leader has access to a data set consisting of tuples of contexts and follower types drawn from a fixed and unknown distribution. Given new contextual information drawn from the same distribution, the leader aims to choose a strategy that maximizes her payoff against a follower whose type is unknown. As in the online setting, this comes with the promise that there exists a function, $h^{*}$, in a class, $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$, that perfectly predicts each follower type given the context. We denote a learning algorithm for distributional learning in structured Stackelberg games as a mapping $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$. Given a context, $\mathbf{z}$, we write the leader’s loss with respect to $h^{*}$, when playing strategy $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ as
$r(\mathbf{z},\hat{\mathbf{x}},f^{(h^{*}(\mathbf{z}))})=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u\left(\mathbf{z},\mathbf{x},b_{f^{(h^{*}(\mathbf{z}))}}(\mathbf{z},\mathbf{x})\right)-u\left(\mathbf{z},\hat{\mathbf{x}},b_{f^{(h^{*}(\mathbf{z}))}}(\mathbf{z},\hat{\mathbf{x}})\right)$
We handle two variants of the notion of sample complexity, which will be central in showing our results. We begin with the classic PAC sample complexity, which we adapt to our setting.
Definition 4.6 (PAC Sample complexity).
@@ 4.2 Learning with Contexts: Problem Setting and Definitions
The PAC sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta)$ of a function class $\pazocal{H}$ with respect to a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, where the infimum is over all possible learning algorithms. With $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$ we denote the minimal integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, every distribution $\mathscr{D}$ on $\pazocal{Z}$, and any true target hypothesis $h^{*}\in\pazocal{H}$, the expected loss of $A$,
$\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[r(\mathbf{z},A(S)(\mathbf{z}),f^{(h^{*}(\mathbf{z}))})\right],$
is at most $\epsilon$ with probability $1-\delta$ over the set $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$, where $\forall i\in[m]\;\mathbf{z}_{i}\sim\mathscr{D}$.
Definition 4.7 (PAC Cut-off sample complexity).
The PAC sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ of a function class $\pazocal{H}$ with respect to a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, where $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ is the minimal integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, every distribution $\mathscr{D}$ on $\pazocal{Z}$, and any true target hypothesis $h^{*}\in\pazocal{H}$, the expected cut-off loss of algorithm $A$,
$\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(S)(\mathbf{z}),f^{(h^{*}(\mathbf{z}))})>\gamma],$
is at most $\epsilon$ with probability $1-\delta$ over the set $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$, where $\forall i\in[m]\;\mathbf{z}_{i}\sim\mathscr{D}$.
For the remainder of this section our results will be with respect to the PAC cut-off sample complexity. Any bound with respect to this benchmark immediately implies a bound on the PAC sample complexity as well. It is not hard to generalize the result of Lemma˜4.8 to our setting, so we omit the proof.
Lemma 4.8 (Lemma 1 in [3]).
@@
4.3 Lower bound
In this section, we introduce the $\gamma$-valued Stackelberg-Natarajan dimension, a new measure of complexity that characterizes a structured Stackelberg game and, as we will see, controls the lower bound of the sample complexity. Indeed, we prove that the finiteness of this measure is a necessary condition for learnability in our setting. This definition is inspired by the classic Natarajan dimension from multiclass prediction [31, 21], and more recently, the $\gamma$-Natarajan dimension from realizable regression [3]. We start with introducing the notion of a $\gamma$-SN-shattered set. Informally, a set of contexts is $\gamma$-SN-shattered if $\pazocal{H}$ can fit complex patterns of follower types on the set. Further, these follower types must be sufficiently different, such that if the type is chosen adversarially among them, the leader incurs loss at least $\gamma$ regardless of her strategy.
Definition 4.9 ($\gamma$-SN-Shattered Set).
Let $\pazocal{H}:[K]^{\pazocal{Z}}$ and $\pazocal{G}$ be a Stackelberg game. We say that a set $S\in\pazocal{Z}^{n}$ is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$ if there exist $g_{0},g_{1}:\pazocal{Z}\to[K]$ such that
$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g_{0}(\mathbf{z}_{i}),g_{1}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})>\gamma\;\;\forall\mathbf{z}_{i}\in S$
and for all $b\in\{0,1\}^{n}$ there exists an $h_{b}\in\pazocal{H}$ such that $h_{b}(\mathbf{z}_{i})=\begin{cases}g_{0}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=0\\ g_{1}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=1\end{cases}$.
Definition 4.10 ($\gamma$-valued Stackelberg-Natarajan Dimension).
The $\gamma$-valued Stackelberg-Natarajan dimension, denoted as $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$, is the cardinality of the largest set that is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
Before stating our main result of this section, we point out that it is important that the two functions, $g_{0}$ and $g_{1}$ be $\gamma$-far from each other in a minmax sense. This allows us to prove a lower bound that holds for any learning algorithm.
Theorem 4.11.
Let $A$ be any learning algorithm and $\epsilon,\delta,\gamma\in(0,1)^{3}$ such that $\delta<\epsilon$, then
@@ 4.3 Lower bound
Let $n_{0}=\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$. Then we know that for any $1\leq n\leq n_{0}$ there exists a set $S_{n}\in\pazocal{Z}^{n}$ and functions $g_{0},g_{1}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that for any set $T\subseteq S$ there exists an $h$ that agrees with $g_{0}$ on all $\mathbf{z}\in T$ and with $g_{1}$ for any $\mathbf{z}\in S\setminus T$. We will show that for a given learning algorithm $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$, there exists a distribution $\mathscr{D}$ on $\pazocal{Z}$ such that the algorithm requires at least $C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$ samples. To do that, we will construct a distribution with support the $n_{0}$-sized $\gamma$-shattered set $S$. We take $\delta<1/15$ and define the distribution $\mathscr{D}$ on $S$ as follows.
@@
$\text{er}_{\mathscr{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma].$
It suffices to establish that there exists $h^{*}\in\pazocal{H}$ such that $\mathbb{P}_{X\sim\mathscr{D}^{m}}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15$. First consider the event
@@
$\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma\land\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n_{0}}\}],$
which trivially lower bounds $\text{er}_{\mathscr{D},\gamma,h}(A(X))$. Also, we consider the following way to choose a true hypothesis $h$. First, we choose vector $b$ uniformly at random from the set $\{0,1\}^{n_{0}}$. We then set $h(\mathbf{z}_{i})=g_{0}(\mathbf{z}_{i})$ for all $b_{i}=0$ and $h(\mathbf{z}_{i})=g_{1}(\mathbf{z}_{i})$ for all $b_{i}=1$. We note that choosing a random $h$ in this way is equivalent to flipping a fair coin for each point in $S$ to determine its label. Since for any point $\mathbf{z}_{i}\in U$ both labels $g_{0}(\mathbf{z}_{i})$ and $g_{1}(\mathbf{z}_{i})$ are consistent with the training set $X$ (by definition of an SN-shattered set), we have that $A(X)$ is independent of the labeling of points in $U$. We also know that if we condition on the point drawn from $\mathscr{D}$ being some unseen point $\mathbf{u}\in U$, we get that for any strategy $A(X)(\mathbf{u})\in\Delta(\pazocal{A})$ of the learner
$\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\geq\begin{cases}1\;\;\text{if}\;h(\mathbf{u})=\arg\max_{j\in\{g_{0}(\mathbf{u}),g_{1}(\mathbf{u})\}}r(\mathbf{u},A(X)(\mathbf{u}),f^{(j)})\\ 0\;\;\text{otherwise}.\end{cases}$
Since we are randomly choosing $h(\mathbf{z})$ to be $g_{0}(\mathbf{z}_{i})$ or $g_{1}(\mathbf{z}_{i})$, we have that the two conditions above happen with probability 1/21/21 / 2 each. Given that there are at least $(d-1)/2$ unseen points, each sampled with probability $\frac{16\epsilon}{d-1}$, and each forcing the learner to make a $\gamma$-error with probability 1/21/21 / 2, we get that
$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;B\right]>\frac{1}{2}\cdot\frac{16\epsilon}{d-1}\cdot\frac{(d-1)}{2}=4\epsilon.$
@@
$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\right]$ $=\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;B\right]\cdot\mathbb{P}[B]$
@@
So there must exists some $h^{*}$ such that $\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))\right]>2\epsilon$. We take $h^{*}$ as the target concept and show that $A$ is likely to produce high error rate. We know that
$\mathbb{P}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h}(A(X))>\epsilon\right]$
$+(1-\mathbb{P}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\leq\epsilon\right]>2\epsilon$
Given that for any algorithm, $A$, it holds that $\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\leq\mathbb{P}[\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}]=16\epsilon$ we get
$\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h}(A(X))>\epsilon\right]\leq 16\epsilon\;\;\text{for any}\;\;h$
@@
$\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot 16\epsilon+(1-\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\epsilon>2\epsilon$
$\iff\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15.$
@@
4.4 Upper bound
In this section, we introduce the $\gamma$-valued Stackelberg-Graph dimension, which, as we will see, is a sufficient condition for distributional learning in structured Stackelberg games. It is also a natural relaxation of the $\gamma$-Natarajan dimension and an adaptation of similar dimensions from classification [31, 21] and regression [3] to our setting.
Definition 4.12 ($\gamma$-SG-Shattered Set).
Let $\pazocal{H}:[K]^{\pazocal{Z}}$ and $\pazocal{G}$ be a Stackelberg game. We say that a sample $S\in\pazocal{Z}^{n}$ is shattered by $\pazocal{H}$ with respect to $\pazocal{G}$ if there exists a function $g:\pazocal{Z}\to[K]$ such that for all $b\in\{0,1\}^{n}$ there exists an $h_{b}\in\pazocal{H}$ such that:
  • $h_{b}(\mathbf{z}_{i})=g(i)\;\;\forall i\in[n]$ such that $b_{i}=0$, and
  • $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g(\mathbf{z}_{i}),h_{b}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})\geq\gamma\;\;\forall i\in[n]$ such that $b_{i}=1$.
Definition 4.13 ($\gamma$-valued Stackelberg-Graph Dimension).
The $\gamma$-valued Stackelberg-Graph dimension, denoted as $\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$ is the cardinality of the largest set that is $\gamma$-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
We are not ready to show our main result in this section, which shows that the $\gamma$-Graph dimension upper bounds the PAC sample complexity in our setting. We design the following simple, improper algorithm that achieves the corresponding bound.
Algorithm 2 Stackelberg Distributional Learner ($\mathfrak{L}^{*}$)
1:input: sample $S=\{(\mathbf{z}_{1},f_{1}),\dots,(\mathbf{z}_{n},f_{n})\}$ that is realizable by $\pazocal{H}$ and test point $\mathbf{z}\in\pazocal{Z}$
2:Consider the subset of hypotheses consistent with the sample, $\pazocal{H}_{c}=\pazocal{H}|_{S}$.
3:Consider the set of followers predicted by $\pazocal{H}_{c}$ on test point $\mathbf{z}$, $F=\pazocal{H}_{c}(\mathbf{z})$
4:Compute and return $\mathbf{x}^{*}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in F}r(\mathbf{z},\mathbf{x},f^{(i)})$
5:end
Theorem 4.14.
For any class $\pazocal{H}:[K]^{\pazocal{Z}}$, any Stackelberg game, $\pazocal{G}$, and for any $\epsilon,\delta,\gamma\in(0,1)$, there exists some constant $C_{1}$ such that Algorithm˜2 achieves:
@@ 4.4 Upper bound
$\text{er}_{\mathscr{D},\gamma}(h):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\right]$
@@
$Q_{\epsilon,\gamma}=\{M\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{M}(h)=0,\text{er}_{\mathscr{D},\gamma}(h)>\epsilon\}$
Now using a classic symmetrization argument [35, 3], we can upper bound the probability of $Q_{\epsilon,\gamma}$ with that of the following set

### v3

@@ B.2 Alternate definition of the Stackelberg Littlestone dimension
It suffices to show that for any finite depth $d$ edge-weighted tree with all root-to-leaf paths having cumulative weight at least $\kappa$ can be shattered under Definition˜B.4 if and only if there exists a tree with root node weight $\kappa$ that can be shattered under Definition˜3.7. We proceed by induction on the depth.
(Base Case) Consider a depth-1 tree, $\pazocal{T}_{1}$, with root node $\mathbf{z}_{\varnothing}$. In the node-weight formulation, $\rho_{\varnothing}=\inf_{\mathbf{x}\in\Delta(\mathcal{A})}\max_{i\in\mathcal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$. In the edge-weight formulation, the valid edge-weight assignments are $\mathcal{N}_{\mathbf{z}_{\varnothing}}=\Bigl\{N=\{\nu_{\varnothing}^{(i)}\}_{i\in\mathcal{H}(\mathbf{z}_{\varnothing})}:\;\forall\mathbf{x}\;\exists i\;r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\nu_{\varnothing}^{(i)}\Bigr\}$. Their minimal path weight is $\max_{N\in\mathcal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\mathcal{H}(\mathbf{z}_{\varnothing})}\nu_{\varnothing}^{(i)}$.
We will first show that $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}$. Fix some $N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$. Then for every $\mathbf{x}\in\Delta(\pazocal{A})$, $\exists i\in\pazocal{H}(\mathbf{z}_{\varnothing})$ such that $r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\nu_{\varnothing}^{(i)}$. This means that
Taking the infimum over $\mathbf{x}$ and the maximum over $N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$ yields the inequality.
For the reverse inequality, let $\mathbf{x}^{*}\in\arg\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$ and $N^{*}=\{\nu_{\varnothing}^{(i)}:=\arg\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x}^{*},f^{(i)})\}_{i\in[K]}$. We argue that $N^{*}\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$, i.e., that
$\forall\mathbf{x}\in\Delta(\pazocal{A})\;\;\exists i\in\pazocal{H}(\mathbf{z}_{\varnothing})\;\;\text{such that}\;\;r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\arg\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z},\mathbf{x}^{*},f^{(i)})$
which holds since $\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$.
(Inductive Step) Assume that the equivalence holds for all trees of depth d1d-1. Consider such a depth-$d$ tree, $\pazocal{T}_{d}$, with root node weight $\mathbf{z}_{\varnothing}$ and children $\mathbf{z}_{\varnothing i}$ for $i\in\pazocal{H}(\mathbf{z}_{\varnothing})$. We want to show that the root node weight $\rho_{\varnothing}$ defined as in Definition˜3.7 is equal to the weight of the minimum path in the tree, where the edge weights are defined as in Definition˜B.4. Concretely, we want to show that
$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl[r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(j)})+\rho_{\varnothing j}\biggr]$ $=\sup_{\pazocal{N}_{d}}\inf_{s\in S_{d}}\Biggl[\sum_{i=0}^{d}\nu_{s[:i]}^{(s[i+1])}\Biggr],$
$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl[r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(j)})+\rho_{\varnothing j}\biggr]$ $=\max_{\pazocal{N}_{0}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\Biggl[\nu_{\varnothing}^{(s[1])}+\max_{\pazocal{N}_{d-1}}\inf_{s\in S_{d-1}}\sum_{i=1}^{d}\nu_{s[:i]}^{(s[i+1])}\Biggr]$
We know that for any $j\in\pazocal{H}(\mathbf{z}_{\varnothing})$ and $s\in S_{d}$ such that $s[1]=j$, $\rho_{\varnothing j}=\max_{\pazocal{N}_{d-1}}\inf_{s\in S_{d-1}}\sum_{i=1}^{d}\nu_{s[:i]}^{(s[i+1])}$ by the inductive hypothesis. Therefore, using an argument similar to the base case, the equation above holds. ∎
Appendix C Details and proofs from Section˜4
In our work, we consider two variants of sample complexity, PAC cut-off sample complexity (Definition˜4.1) and PAC sample complexity defined as follows.
Definition C.1 (PAC sample complexity).
@@ Appendix C Details and proofs from Section˜4
The PAC sample complexity of a function class $\pazocal{H}$ w.r.t. a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, where the infimum is over all learning algorithms. $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$ is the smallest integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, every distribution $\mathcal{D}$ on $\pazocal{Z}$, and any true hypothesis $h^{*}\in\pazocal{H}$, the expected loss of $A$, $\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathcal{D}}\left[r(\mathbf{z},A(\mathbf{z};S),f^{(h^{*}(\mathbf{z}))})\right]$, is at most $\epsilon$ with probability at least $1-\delta$, where $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$ and $\mathbf{z}_{i}\sim\mathcal{D}$.
According to the following lemma, any bound with respect to the PAC cut-off complexity immediately implies a bound on the PAC sample complexity as well.
Lemma C.2 (Lemma 1 in [Attias23:regression]).
@@
See 4.4
@@
Let $n_{0}=\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$. Then we know that for any $1\leq n\leq n_{0}$ there exists a set $S_{n}\in\pazocal{Z}^{n}$ and functions $g_{0},g_{1}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that for any set $T\subseteq S$ there exists an $h$ that agrees with $g_{0}$ on all $\mathbf{z}\in T$ and with $g_{1}$ for any $\mathbf{z}\in S\setminus T$. We will show that for a given learning algorithm $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$, there exists a distribution $\mathcal{D}$ on $\pazocal{Z}$ such that the algorithm requires at least $C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$ samples. To do that, we will construct a distribution with support the $n_{0}$-sized $\gamma$-shattered set $S$. We take $\delta<1/15$ and define the distribution $\mathcal{D}$ on $S$ as follows.
@@
$\text{er}_{\mathcal{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathcal{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma].$
It suffices to establish that there exists $h^{*}\in\pazocal{H}$ such that $\mathbb{P}_{X\sim\mathcal{D}^{m}}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15$. First consider the event
@@
$\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))$ $:=\mathbb{P}_{\mathbf{z}\sim\mathcal{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})$
$>\gamma\land\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n_{0}}\}],$
which lower bounds $\text{er}_{\mathcal{D},\gamma,h}(A(X))$. Also, we consider the following way to choose a true hypothesis $h$. First, we choose vector $b$ uniformly at random from the set $\{0,1\}^{n_{0}}$. We then set $h(\mathbf{z}_{i})=g_{0}(\mathbf{z}_{i})$ for all $b_{i}=0$ and $h(\mathbf{z}_{i})=g_{1}(\mathbf{z}_{i})$ for all $b_{i}=1$. We note that choosing a random $h$ in this way is equivalent to flipping a fair coin for each point in $S$ to determine its label. Since for any point $\mathbf{z}_{i}\in U$ both labels $g_{0}(\mathbf{z}_{i})$ and $g_{1}(\mathbf{z}_{i})$ are consistent with the training set $X$ (by definition of an SN-shattered set), we have that $A(X)$ is independent of the labeling of points in $U$. We also know that if we condition on the point drawn from $\mathcal{D}$ being some unseen point $\mathbf{u}\in U$, we get that for any strategy $A(X)(\mathbf{u})\in\Delta(\pazocal{A})$ of the learner $\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\geq 1$ if $h(\mathbf{u})=\arg\max_{j\in\{g_{0}(\mathbf{u}),g_{1}(\mathbf{u})\}}r(\mathbf{u},A(X)(\mathbf{u}),f^{(j)})$ and 0 otherwise.
Since we are randomly choosing $h(\mathbf{z})$ to be $g_{0}(\mathbf{z}_{i})$ or $g_{1}(\mathbf{z}_{i})$, we have that the two conditions above happen with probability $1/2$ each. Given that there are at least (d1)/2(d-1)/2 unseen points, each sampled with probability $\frac{16\epsilon}{d-1}$, and each forcing the learner to make a $\gamma$-error with probability $1/2$, we get that
$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\;|\;B\right]>\frac{1}{2}\cdot\frac{16\epsilon}{d-1}\cdot\frac{(d-1)}{2}=4\epsilon.$
@@
$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\right]$ $=\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\;|\;B\right]\cdot\mathbb{P}[B]$
@@
So there must exists some $h^{*}$ such that $\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h^{*}}(A(X))\right]>2\epsilon$. We take $h^{*}$ as the target concept and show that $A$ is likely to produce high error rate. We know that
$\mathbb{P}_{X}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathcal{D},\gamma,h}(A(X))>\epsilon\right]$
$+(1-\mathbb{P}_{X}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))\leq\epsilon\right]>2\epsilon$
Given that for any algorithm, $A$, it holds that $\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\leq\mathbb{P}[\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}]=16\epsilon$ we get
$\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\;|\;\text{er}_{\mathcal{D},\gamma,h}(A(X))>\epsilon\right]\leq 16\epsilon\;\;\text{for any}\;\;h$
@@
$\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot 16\epsilon+(1-\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\epsilon>2\epsilon$
$\iff\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15.$
@@
See 4.7
@@
$\text{er}_{\mathcal{D},\gamma}(h):=\mathbb{P}_{\mathbf{z}\sim\mathcal{D}}\left[\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\right]$
@@
$Q_{\epsilon,\gamma}=\{M\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{M}(h)=0,\text{er}_{\mathcal{D},\gamma}(h)>\epsilon\}$
Now using a classic symmetrization argument [uml, Attias23:regression], we can upper bound the probability of $Q_{\epsilon,\gamma}$ with that of the following set